A committee of 6 can be formed from a group of 5 sophomores, 7 juniors, and 8 senior. In how many ways can these be done if:



A. there are 2 members from each class on the committees



B. there shall be exactly 3 seniors, 2 juniors, and 1 sophomore on the committee

1
Expert's answer
2022-02-25T11:26:11-0500

Let "A" denote the event ''sophomore is chosen'', "B" denote the event ''junior is chosen'', "C" denote the event ''senior is chosen'',

There are "\dbinom{5+7+8}{6}=38760" possible outcomes.

A.


"P(2A, 2B, 2C)=\dfrac{\dbinom{5}{2}\dbinom{7}{2}\dbinom{8}{2}}{\dbinom{20}{6}}"

"=\dfrac{10(21)(28)}{38760}=\dfrac{49}{323}=0.1517"

B.


"P(1A, 2B, 3C)=\dfrac{\dbinom{5}{1}\dbinom{7}{2}\dbinom{8}{3}}{\dbinom{20}{6}}"

"=\dfrac{5(21)(56)}{38760}=\dfrac{49}{323}=0.1517"